# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 7.1. Rank one examples
The examples 7.1-7.2-7.3-7.4 below, as well as Example 7.9 in the following section, illustrate webs of $5$-branes with $7$-branes engineering the rank one $E_{0}$ $5$d SCFT obtained by $M$-theory on local $\mathbb{P}^{2}$.
@@
Let $\overline{P}$ be the lattice polygon shown on the left of Figure 7.1. The toric surface $(\overline{Y},\overline{D})$ with momentum polytope $\overline{P}$ is singular and has an $A_{2}$ singularity at each of the three 0-dimensional strata of $\overline{D}$. As discussed in §2.2.3, the polyhedral decomposition $\mathcal{\overline{P}}$ of $\overline{P}$ in the middle of Figure 7.1 determines a one-parameter toric degeneration of $\overline{Y}$ into the union of three copies of $\mathbb{P}^{2}$, corresponding to the three triangles forming $\overline{P}$. According to §2.3, the mirror $\mathcal{\overline{X}}$ to this one-parameter degeneration of $\overline{Y}$ is the toric Calabi–Yau 3-fold whose fan is the cone over $(\overline{P},\mathcal{\overline{P}})$. In other words, the momentum polytope image of $\mathcal{\overline{X}}$ is the cone over the dual web illustrated in blue on the right Figure 7.1. In this situation, we have $\mathcal{\overline{X}}=K_{\mathbb{P}^{2}}$. Moreover, the corresponding map $\mathcal{\overline{X}}=K_{\mathbb{P}^{2}}\to\mathbb{C}$ is determined by the section $s=xyz\in H^{0}(\mathbb{P}^{2},-K_{\mathbb{P}^{2}})$, where $x,y,z$ denote the homogeneous coordinates of $\mathbb{P}^{2}$. In particular, the intersection of $\mathbb{P}^{2}$ with the other irreducible components of the central fiber $\overline{\mathcal{X}}_{0}$ is the union of the three lines forming the toric boundary of $\mathbb{P}^{2}$ and defined by the equation $s=0$. Contracting the zero-section $\mathbb{P}^{2}\subset\mathcal{X}$ to a point, we obtain an affine toric canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$, which is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with three irreducible components.
[…]
Figure 7.1. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.2. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by one interior blow-up from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.2 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to a union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 7.2 is obtained from $\mathbb{P}^{2}$ by smoothing one of the corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by the union of a line and a conic. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with two irreducible components.
[…]
Figure 7.2. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.3. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by two interior blow-ups from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.3 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to a non-normal surface whose normalization is isomorphic to $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 7.3 is obtained from $\mathbb{P}^{2}$ by smoothing two of the corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by a nodal cubic curve. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with a unique non-normal, self-intersecting, irreducible component.
[…]
Figure 7.3. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.4. The corresponding log Calabi–Yau surface with line bundle $(Y,D,L)$ is obtained by three interior blow-ups from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.4 and corresponds to a one-parameter degeneration of $(Y,D)$. In this case, we have $L^{2}=0$, and contracting all curves having zero-intersection with $L$ produces a map $c:Y\rightarrow\mathbb{C}$ which is an elliptic fibration. As described in §6.2, the mirror $\mathcal{X}$ is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 7.4 is obtained from $\mathbb{P}^{2}$ by smoothing the three corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by a smooth cubic curve. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a simple elliptic singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$.
[…]
Figure 7.4. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.
The following example illustrates a web of $5$-branes with $7$-branes, describing the $\widetilde{E}_{1}$ $5$d SCFT obtained by $M$-theory on local $\mathbb{F}_{0}$.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.5, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by a blow-up of $\overline{Y}$ along a smooth point on the component of the boundary divisor corresponding to the bottom edge of $\overline{P}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.5 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of three copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of the toric Calabi–Yau 3-fold $\overline{\mathcal{X}}=K_{\mathbb{F}_{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 7.5 is obtained from $\mathbb{F}_{2}$ by smoothing one corner of its toric boundary. This compact component is isomorphic to $\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, with an intersection with the other irreducible components of $\mathcal{X}_{0}$ given by a non-toric divisor with three irreducible components. In particular, we have $\mathcal{X}=K_{\mathbb{F}_{0}}$, but with a non-toric morphism $\mathcal{X}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with three irreducible components.
[…]
Figure 7.5. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $\widetilde{E}_{1}$ 5d SCFT.
The following example illustrates a web of $5$-branes with $7$-branes, describing the $E_{1}$ $5$d SCFT obtained by $M$-theory on local $\mathbb{F}_{1}$.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.6, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by a blow-up of $\overline{Y}$ along two smooth points on two components of the boundary divisor corresponding to the most left, and bottom edges of $\overline{P}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.6 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 7.6 that the central fiber $\mathcal{X}_{0}$ consists of three irreducible component. One irreducible component is compact and isomorphic to $\mathbb{P}^{2}$. The other two components are non-compact, one being toric and the other being non-toric and containing an interior $(-1)$-curve. Flopping this curve, we obtain the Calabi–Yau $3$-fold $K_{\mathbb{F}_{1}}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with two irreducible components.
[…]
Figure 7.6. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{1}$ 5d SCFT.

### v2

@@ 7.1. Rank one examples
The examples 7.1-7.2-7.3-7.4 below, as well as Example 7.9 in the following section, illustrate webs of $5$-branes with $7$-branes engineering the rank one $E_{0}$ 55d SCFT obtained by $M$-theory on local $\mathbb{P}^{2}$.
@@
Let $\overline{P}$ be the lattice polygon shown on the left of Figure 28. The toric surface $(\overline{Y},\overline{D})$ with momentum polytope $\overline{P}$ is singular and has an $A_{2}$ singularity at each of the three 0-dimensional strata of $\overline{D}$. As discussed in §2.2.3, the polyhedral decomposition $\mathcal{\overline{P}}$ of $\overline{P}$ in the middle of Figure 28 determines a one-parameter toric degeneration of $\overline{Y}$ into the union of three copies of $\mathbb{P}^{2}$, corresponding to the three triangles forming $\overline{P}$. According to §2.3, the mirror $\mathcal{\overline{X}}$ to this one-parameter degeneration of $\overline{Y}$ is the toric Calabi–Yau 3-fold whose fan is the cone over $(\overline{P},\mathcal{\overline{P}})$. In other words, the momentum polytope image of $\mathcal{\overline{X}}$ is the cone over the dual web illustrated in blue on the right Figure 28. In this situation, we have $\mathcal{\overline{X}}=K_{\mathbb{P}^{2}}$. Moreover, the corresponding map $\mathcal{\overline{X}}=K_{\mathbb{P}^{2}}\to\mathbb{C}$ is determined by the section $s=xyz\in H^{0}(\mathbb{P}^{2},-K_{\mathbb{P}^{2}})$, where $x,y,z$ denote the homogeneous coordinates of $\mathbb{P}^{2}$. In particular, the intersection of $\mathbb{P}^{2}$ with the other irreducible components of the central fiber $\overline{\mathcal{X}}_{0}$ is the union of the three lines forming the toric boundary of $\mathbb{P}^{2}$ and defined by the equation $s=0$. Contracting the zero-section $\mathbb{P}^{2}\subset\mathcal{X}$ to a point, we obtain an affine toric canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$, which is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with three irreducible components.
Figure 28. On the left, the momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$. In the middle, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$. On the right, the associated dual web of $5$-branes engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 29. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by one interior blow-up from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 29 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to a union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 29 is obtained from $\mathbb{P}^{2}$ by smoothing one of the corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by the union of a line and a conic. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with two irreducible components.
Figure 29. On the left, the momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$. In the middle, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$. On the right, the associated dual web of $5$-branes engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 30. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by two interior blow-ups from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 30 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to a non-normal surface whose normalization is isomorphic to $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 30 is obtained from $\mathbb{P}^{2}$ by smoothing two of the corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by a nodal cubic curve. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with a unique non-normal, self-intersecting, irreducible component.
Figure 30. On the left, the momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$. In the middle, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$. On the right, the associated dual web of $5$-branes engineering the rank one $E_{0}$ 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 31. The corresponding log Calabi–Yau surface with line bundle $(Y,D,L)$ is obtained by three interior blow-ups from the toric surface $(\overline{Y},\overline{D})$ as in Example 7.1. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 31 and corresponds to a one-parameter degeneration of $(Y,D)$. In this case, we have $L^{2}=0$, and contracting all curves having zero-intersection with $L$ produces a map $c:Y\rightarrow\mathbb{C}$ which is an elliptic fibration. As described in §6.2, the mirror $\mathcal{X}$ is a non-toric deformation of $\overline{\mathcal{X}}=K_{\mathbb{P}^{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 31 is obtained from $\mathbb{P}^{2}$ by smoothing the three corners of its toric boundary. This compact component is still isomorphic to $\mathbb{P}^{2}$, but its intersection with the other irreducible components of $\mathcal{X}_{0}$ is the non-toric divisor defined by a smooth cubic curve. It follows that we still have $\mathcal{X}=K_{\mathbb{P}^{2}}$, but the map $\mathcal{X}\rightarrow\mathbb{C}$ is a non-toric deformation of the toric morphism $\overline{\mathcal{X}}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a simple elliptic singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$.
Figure 31. On the left, the momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$. In the middle, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$. On the right, the associated dual web of $5$-branes engineering the rank one $E_{0}$ 5d SCFT.
The following example illustrates a web of $5$-branes with $7$-branes, describing the $\widetilde{E}_{1}$ 55d SCFT obtained by $M$-theory on local $\mathbb{F}_{0}$.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 32, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by a blow-up of $\overline{Y}$ along a smooth point on the component of the boundary divisor corresponding to the bottom edge of $\overline{P}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 32 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of three copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of the toric Calabi–Yau 3-fold $\overline{\mathcal{X}}=K_{\mathbb{F}_{2}}$. In particular, the compact component of $\mathcal{X}_{0}$ corresponding to the bounded cell in the dual web on the right of Figure 32 is obtained from $\mathbb{F}_{2}$ by smoothing one corner of its toric boundary. This compact component is isomorphic to $\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, with an intersection with the other irreducible components of $\mathcal{X}_{0}$ given by a non-toric divisor with three irreducible components. In particular, we have $\mathcal{X}=K_{\mathbb{F}_{0}}$, but with a non-toric morphism $\mathcal{X}\rightarrow\mathbb{C}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with three irreducible components.
Figure 32. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $\widetilde{E}_{1}$ 5d SCFT.
The following example illustrates a web of $5$-branes with $7$-branes, describing the $E_{1}$ 55d SCFT obtained by $M$-theory on local $\mathbb{F}_{1}$.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 33, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by a blow-up of $\overline{Y}$ along two smooth points on two components of the boundary divisor corresponding to the most left, and bottom edges of $\overline{P}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 33 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 33 that the central fiber $\mathcal{X}_{0}$ consists of three irreducible component. One irreducible component is compact and isomorphic to $\mathbb{P}^{2}$. The other two components are non-compact, one being toric and the other being non-toric and containing an interior (1)(-1)-curve. Flopping this curve, we obtain the Calabi–Yau $3$-fold $K_{\mathbb{F}_{1}}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a degenerate cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$ with two irreducible components.
Figure 33. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{1}$ 5d SCFT.